# HallE-Switch: Rethinking and Controlling Object Existence Hallucinations in Large Vision Language Models for Detailed Caption

### v1

@@ A.5 Theoretical Explanations
Inspired by the Theorem, we propose a linear transform in word embedding space for LVLM. Let be the finetuned LLM, with switch layer/projector denoted as , we replace each word embedding with , making the new language model as HallE-Switch’s language decoder. is adjustable from -1 to 1. We assign and fit contextual only data; assign with the contextual and parametric combined data. After finetuning the HallE-Switch, the user only needs to specify a switch value and do normal vision language task like image captions. We use maximal likelihood as the training objective.
@@
The design of LM-Switch maintains a linearity guarantee, with proof of the switch model’s distribution is close to a linear interpolation. Let be the maximum eigen-value of W. When varying s value,
distribution of the switch model is close a linear interpolation of and , meaning that the model distribution changes linearly. Therefore, our method can take any between 1 and -1.
![[Uncaptioned image]](extracted/5143148/figures/000000146988.jpg)
![[Uncaptioned image]](extracted/5143148/figures/000000157923.jpg)
![[Uncaptioned image]](extracted/5143148/figures/000000192920.jpg)

### v3

@@ A.11 Theoretical Explanations
Inspired by the Theorem, we propose a linear transform in word embedding space for LMM. Let be the finetuned LLM, with control layer/projector denoted as , we replace each word embedding with , making the new language model as HallE-Control’s language decoder. is adjustable from -1 to 1. We assign and fit contextual only data; assign with the contextual and parametric combined data. After finetuning the HallE-Control, the user only needs to specify a control value and do normal vision language task like image captions. We use maximal likelihood as the training objective.
@@
The design of LM-Switch maintains a linearity guarantee, with proof of the control model’s distribution is close to a linear interpolation. Let be the maximum eigen-value of W. When varying s value,
distribution of the control model is close a linear interpolation of and , meaning that the model distribution changes linearly. Therefore, our method can take any between 1 and -1.
For HallE-Control, the core idea is: Assuming LLM is good enough to represent an equivalent distribution with HMM; there exist an matrix W, so that after transferring word embedding to , the LLM’s originally simulate the text distribution starting with initial state will turn to be equivalent to a distribution starting with initial state .
According to experiments in Section 4.1, it reveals language models can distinguish inference object and observed objects by adding special tokens. Because, decoder based language models generate sequences in an auto-regressive way. We followed the proof in Han et al. (2023): if we assume O as our observation space which means a set of m observations. We assume as a contextual vector, as word embedding. We can represent word logits as E. In attention mechanic, it will pass through a softmax operator to get the distribution over words, We use the same assumption in LM-Switch method to assume a linear formulation and let the conditional probability in language model . We can get the full probability by using chain-rule: . We assume our LLMs are good enough to represent an equivalent distribution with HMM and full column-rank for E, p(o).
Our conclusion is: Applying a linear transformation on word embedding space is equivalent to a shift from one initial condition to another. This is the reason that we want to shift a language model with distribution produce higher inference imagination conditional probability to a lower one.
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000146988.jpg)
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000157923.jpg)
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000192920.jpg)